# Stability-Aware Training of Neural Network Interatomic Potentials with Differentiable Boltzmann Estimators

### v1

@@ 3.3 Ac-Ala3-NHMe Tetrapeptide
Ac-Ala3-NHMe (chemical formula $C_{12}H_{22}N_{4}O_{4}$) is a tetrapeptide from the MD22 dataset [undefp], a challenging benchmark consisting of reference energy and forces computed at the PBE+MBD [undefax, undefaan] level of theory for considerably larger molecules than those in the MD17 dataset. Relative to MD17 molecules, Ac-Ala3-NHMe poses unique challenges for atomistic simulations due to its larger size and flexibility [undefae]. As a result, more expressive NNIPs—often incorporating E(3) equivariance [undefz]—are required to accurately model the underlying potential energy surface.
Problem setup.
In our experiments, we use a NequIP [undefd] model due to its promising accuracy and data efficiency on challenging atomistic simulation tasks [undefaj, undefar]. We pre-train a NequIP model on the QM energy and forces matching objective using 14,890 (or 25%) of available structures from the training dataset. After convergence, we begin StABlE Training with the global Boltzmann estimator, using 48 parallel replicas at a temperature of 500K. As with aspirin (Sections 3.1 - 3.2), we detect instability using the maximum bond length deviation criterion, and choose the distribution of interatomic distances, $h(r)$, as our training observable.
Results.
The StABlE-trained NequIP model gives an improvement in stability relative to a baseline model trained only on QM reference data (Figure 5a). Additionally, the StABlE-trained model closely recovers the true $h(r)$, while the baseline NNIP produces a highly noisy $h(r)$. This is because of insufficient sampling, due to a lower stable simulation time (Figure 5b). The baseline NNIP also underestimates the latter portion of the $h(r)$ distribution, indicating poor modelling of long-range interactions. Figure 5c shows selected structures throughout the course of a single learning phase of StABlE Training, demonstrating a progression from very unstable to stable structures.
@@ 3.4 Water
Figure 5: Ac-Ala3-NHMe tetrapeptide simulation with StABlE Training. Results of StABlE Training a NequIP  [undefd] NNIP on the 42-atom Ac-Ala3-NHMe tetrapeptide from the MD22 [undefp] dataset at 500 Kelvin. (a) Fraction of unstable replicas as a function of simulation time for a NequIP NNIP trained on 1,000 energy and forces reference data points (blue). Applying StABlE Training yields a model which can simulate more MD replicas stably over time (orange). (b) Distribution of interatomic distances ($h(r)$) from an ab-initio molecular dynamics (AIMD) simulation (black), a NequIP model trained on 14,890 QM data points (blue), and a NequIP model trained on 14,890 QM datapoints, followed by our StABlE Training procedure (orange). The StABlE-trained NequIP model much more closely recovers the true $h(r)$, while the $h(r)$ produced by the model trained only on QM reference data is noisy and inaccurate due to insufficient sampling time. (c) Ac-Ala3-NHMe structures sampled over epochs of a single learning phase of StABlE Training. There is a clear progression as very unstable configurations become stable.
As a final evaluation of StABlE Training, we consider liquid-phase water. Water has historically posed unique challenges for atomistic simulations due to the presence of long-range interactions, proton disorder, and nuclear quantum effects [undefo, undefay, undefaq]. Additionally, estimating transport properties, such as the diffusivity coefficient, requires long, continuous trajectories to minimize statistical error. We use the dataset produced in [undefv], which contains reference all-atom simulations of 64 water molecules using the flexible Extended Simple Point Charge model [undefaat] at 300K and 1 atm. For this system, we consider a GemNet-T model [undefw]. GemNet-T displays a failure mode in which unphysical configurations (e.g., bond stretching) first arise in highly localized ”pockets” of 1-2 molecules (Figure 6a), which then gradually cascade to the rest of the simulation domain [undefv]. Such configurations do not noticeably affect global system observables such as the radial distribution function. Therefore, in this setting it is more appropriate to use our localized version of the Boltzmann gradient estimator (Equation 8) in order to guide the optimization process towards specific spatial domains of instability.
Problem setup.
We pre-train a GemNet-T model to the energy and forces matching objective using a subset of 1,000 reference structures from the training dataset. After convergence, we begin StABlE Training with the localized Boltzmann estimator, simulating 10 parallel replicas at a temperature of 300K. We use a minimum intermolecular distance criterion (see Supplementary Section A.5) to detect instabilities. To estimate gradients for optimization, from each global state we extract all possible local neighborhoods containing a single water molecule (see Supplementary Section A.6 for more details on the local neighborhood sampling scheme). The most prevalent failure mode in this system is unphysical bond stretching, which eventually leads to collisions with nearby water molecules. Motivated by this observation, we use the mean hydrogen-oxygen bond length as our training observable. We perform a single cycle of simulation and learning.
Results.
We evaluate the performance of the StABlE-Trained model by performing 10 parallel MD simulations at 300K, starting from held-out initial conditions. We achieve significant improvements in stability and accuracy using our training approach. We nearly double the stable simulation time relative to a baseline NNIP model trained only on QM reference data (Figure 6b). Compared to a reference value of $2.3\times 10^{-9}\frac{m^{2}}{s}$, the mean diffusivity produced by StABlE-trained model simulations is $1.6\times 10^{-9}\frac{m^{2}}{s}$, while the mean diffusivity produced by baseline model simulations is $5.0\times 10^{-9}\frac{m^{2}}{s}$ (Figure 6c). This corresponds to an 87% improvement in the accuracy of diffusivity coefficient estimation. We note that the StABlE-trained model systemically underestimates the reference diffusivity, while the baseline model overestimates the value. We hypothesize that in this particular case, StABlE Training has a retarding effect on system dynamics. As the StABlE optimization process assigns high probability to microstates closely matching the equilibrium bond length, small fluctuations about this value may be excessively penalized, inhibiting the motion of molecules in the simulation. Finally, the spikes in the short-range region of the element-conditioned RDFs (Figures 6d-f), which indicate unphysical bond stretching and collision events, are largely removed in StABlE-trained model simulations, while long-range correlations are also better captured. We reiterate that in this experiment, neither the element-conditioned RDFs nor the diffusivity coefficient were explicitly seen during StABlE Training.
@@ Results.
Figure 6: All-Atom Water Simulation with StABlE Training. Results of StABlE Training a GemNet-T  [undefw] NNIP on an all-atom water system with 64 molecules at 300 Kelvin. (a) Example unstable configuration arising in a GemNet-T simulation of water. Instabilities are highly localized and manifest as unphysical bond stretches or intermolecular collisions, motivating the use of our localized Boltzmann estimator (Section 22.1) during training (b) Fraction of unstable replicas as a function of simulation time for a GemNet-T NNIP trained on energy and forces reference data points (blue), followed by StABlE Training (orange). StABlE Training yields a clear stability improvement relative to training on only energy and forces data, more than doubling the median stable simulation time from 23 to 52 picoseconds. (c) Left: Convergence of the diffusivity coefficient over the simulations produced by the reference Extended Simple Point Charge model [undefaat] (black), a GemNet-T model trained on 1,000 energy and forces datapoints (blue), and a GemNet-T model trained on 1,000 energy and forces datapoints followed by StABlE Training (orange). Right: Relative to the reference value of $2.3\times 10^{-9}\frac{m^{2}}{s}$The StABlE-trained model gives a more accurate final estimate of the diffusivity coefficient, with a mean predicted value of $1.6\times 10^{-9}\frac{m^{2}}{s}$ as compared with a value of $5.0\times 10^{-9}\frac{m^{2}}{s}$ produced by the baseline model. (d - f) Element-conditioned (OH, $O$, $H$) radial distribution functions (RDF) produced by simulations from the reference model (black), a GemNet-T model trained on 1,000 energy and forces data points (blue), and a GemNet-T model trained on 1,000 energy and forces data points followed by StABlE Training (orange). The StABlE-trained model produces fewer unphysical collisions and more accurately captures long-range interactions.
4 Discussion
We have introduced StABlE Training, a strategy for training stable and accurate neural network interatomic potentials. We have demonstrated StABlE Training on a flexible organic molecule, a tetrapeptide, and a liquid-phase system, as well as with three modern NNIP architectures. Our training procedure results in NNIPs which remain stable for considerably longer periods in MD simulation compared to models trained only on QM data, and as a result, more accurately reproduce key simulation observables. We also note that we have improved the recovery of observables that the NNIPs were not explicitly trained on. These include dynamical observables such as the diffusivity coefficient of water and velocity autocorrelation function of aspirin. We showed that StABlE-trained models exhibit better generalization to unseen temperatures for the aspirin molecule. Finally, we demonstrated that StABlE-trained models can outperform NNIP models trained on up to 50 times more QM data. We thus expect StABlE Training to be a particularly compelling alternative to traditional QM training in cases where producing comprehensive ab-initio datasets is prohibitively expensive, such as for large system sizes.
Advantages of StABlE Training.
One advantage of StABlE Training is that it can utilize both quantum-mechanical energies and forces and system observables to supervise NNIP training. As reference observables can be easily computed by averaging over existing datasets, the stability, accuracy, and data efficiency gains brought by StABlE Training require no additional reference calculations or data. StABlE Training is thus self-contained and efficient, requiring minimal additional computational expense beyond conventional NNIP training (see Supplementary Section A.6 for details on training times).
StABlE Training can also be flexibly applied across atomistic systems due to its ability to handle a diverse set of failure modes arising in MD simulation, including both global and local instabilities. Specifically, due to our localized Boltzmann estimator, we can target and resolve isolated instabilities found in simulations of large condensed-phase systems. The adaptability of our method is further exemplified by the lack of reliance on a manually specified prior potential, a common requirement of observable-based learning schemes. The pre-training step which precedes StABlE Training effectively learns a QM-informed prior, which is then further refined in the alternating simulation and learning stages.
The flexibility of StABlE Training extends to the neural network parameterization of the potential energy function. The three NNIP architectures we considered in this work vary widely in their expressivity, as measured by the number of total learnable parameters, as well as in the symmetry principles governing their internal feature transformations (see Supplementary Section A.6 for details). The effectiveness of StABlE Training across these models suggests that our approach will remain applicable as NNIP architectures continue to evolve.
Finally, we note that the underlying Boltzmann gradient estimator used to train the NNIPs differentiates through the potential energy of uncorrelated states. It is thus agnostic to the method of sampling from the Boltzmann distribution $P_{\theta}(\Gamma)$, and is readily applicable to sampling schemes other than MD simulation (e.g., Monte Carlo methods). This is related to recent work in machine learning on implicit layers, which enables differentiation of optimization problem solutions without knowledge of the numerical solver used [undef, undefaa, undefaz, undefh, undefav].
@@ 4 Discussion
A limitation of StABlE Training is that training NNIPs to match reference observables lacks the convergence guarantees provided by traditional QM learning in the large data limit. Specifically, the observable-matching objective is under-constrained, as the mapping between a potential energy function and a sparse set of simulation observables is in general non-unique [undefaw]. Although we mitigate this problem by pre-training and regularizing StABlE Training with the conventional energy and forces loss function, we observe that stability improvements resulting from StABlE Training are accompanied by a minor increase in the energy and forces error on a held-out test set, indicating a misalignment between the observable-matching and QM objectives. Adjusting the strength of the QM regularization and the learning rate used for StABlE Training are two primary ways in which to navigate this tradeoff. We note, however, that our observed error increases are typically within the range of DFT error (see Supplementary Section A.10 for details). We also note that StABlE Training is currently incompatible with dynamical observables, due to the use of uncorrelated states to compute the Boltzmann Estimator. Overcoming this limitation is non-trivial, as it requires optimization over long paths, which is a significantly higher-dimensional problem with large-variance gradients [undefi].
@@
In this work, we took observables calculated from high-fidelity simulations as reference values. Many of the observables we considered, such as the radial distribution function, diffusivity coefficient, and equilibrium bond length, are also experimentally measurable. Future work could explore training with experimental observables from multiple thermodynamic conditions (e.g., temperatures and pressures) simultaneously to yield more generalizable and robust potentials. As an independent source of information, experimental observables could also be used to refine NNIPs trained on data obtained from lower levels of theory. Future work could also explore incorporating additional observables into StABlE Training, particularly those which are dynamical. This could restrict the set of learnable functions and address the under-constrained nature of learning with observables. Finally, we note that established techniques, such as active learning [undefaai, undefaao, undefan], are not mutually exclusive with StABlE Training, and could be used in tandem to further improve NNIP stability and robustness.
5 Acknowledgements
This work was initially supported by Laboratory Directed Research and Development (LDRD) funding under Contract Number DE-AC02-05CH11231. It was then supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) program under contract No. DE-AC02-05CH11231. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231. We thank Rasmus Lindrup, Toby Kreiman, Geoffrey Negiar, Muhammad Hasyim, Ritwik Gupta, Aayush Singh, David Limmer, Kranthi Mandadapu, David Prendergast, Bryan McCloskey, and Muratahan Aykol for fruitful discussions and comments on the manuscript.
6 Author Contributions
S.R and A.S.K. jointly conceived and initiated the project. S.R. conceptualized and implemented the Localized Boltzmann Estimator and StABlE Training method, and performed experiments. I.A. trained the baseline models and contributed to the Boltzmann Estimator theory. F.P contributed to a unique derivation and analysis of the Boltzmann Estimator. A.S.K. supervised and guided the project from conception to design of experiments, implementation, theory, and analysis. All authors analyzed results and wrote the paper.
7 Code Availability
Our code will be made available at https://github.com/ASK-Berkeley/StABlE-Training.
Appendix A Supplementary Information
@@ A.1 Derivation of Boltzmann Estimator
We provide a full derivation of our Boltzmann estimator, which we use to train our NNIP as part of StABlE Training. Consider a vector-valued observable $g(\Gamma)$ of a state $\Gamma$, and a reference value of the observable $g_{\text{ref}}$. Training a NNIP $U_{\theta}$ to match $g_{\text{ref}}$ requires minimizing the loss function,
@@
where $P_{\theta}(\Gamma)$ is the equilibrium distribution induced by the NNIP $U_{\theta}$. This requires computing the gradient, $\nabla_{\theta}\mathcal{L}_{\text{obs}}$, which can be decomposed via the chain rule as follows:
@@
We derive the $N$-sample estimator, presented in Equation LABEL:eq:estimator, of the Jacobian, ${\partial\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)}[g(\Gamma)]}/{\partial\theta}$. The estimator is repeated below for convenience.
$\mathcal{E}(\Gamma^{(1)},\ldots,\Gamma^{(N)})\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{N}{k_{B}T(N-1)}\left[\hat{\mathbb{E}}\left[g(\Gamma)\right]\hat{\mathbb{E}}\left[{\nabla_{\theta}U_{\theta}(\Gamma)}\right]^{\top}-\hat{\mathbb{E}}\left[{g(\Gamma)\nabla_{\theta}U_{\theta}(\Gamma)}^{\top}\right]\right],$
where $\hat{\mathbb{E}}$ denotes sample averages. This is an unbiased estimator, that is, $\mathbb{E}_{(\Gamma^{(1)},\ldots,\Gamma^{(N)}\sim P_{\theta}(\Gamma)}[\mathcal{E}(\Gamma^{(1)},\ldots,\Gamma^{(N)})]=\frac{\partial\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)}[g(\Gamma)]}{\partial\theta}$.
@@
Given $P_{\theta}(\Gamma)\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{\exp\left(-\frac{1}{k_{B}T}\mathcal{H}_{\theta}(\Gamma)\right)}{C(\theta)}$, where $\mathcal{H}(\Gamma)=\sum_{i=1}^{N}\frac{p_{i}^{2}}{2m_{i}}+U_{\theta}(r)$, and $C(\theta)\stackrel{{\scriptstyle\text{def}}}{{=}}\int\exp(-\frac{1}{k_{B}T}\mathcal{H}_{\theta}(\Gamma^{\prime}))\mathop{}\!\mathrm{d}\Gamma^{\prime}$ is the partition function, we have,
@@
Expanding via the chain rule for Jacobians, and noting that $\nabla_{\theta}\mathcal{H}_{\theta}(\Gamma)=\nabla_{\theta}U_{\theta}(r)$ (we will write $U_{\theta}(\Gamma)$ for convenience), we get,
$=\int g(\Gamma^{\prime})\frac{-\frac{1}{k_{B}T}\exp\left(-\frac{1}{k_{B}T}\mathcal{H}_{\theta}(\Gamma^{\prime})\right)\nabla_{\theta}U_{\theta}(\Gamma^{\prime})^{\top}C(\theta)-\nabla_{\theta}C(\theta)^{\top}\exp\left(-\frac{1}{k_{B}T}\mathcal{H}_{\theta}(\Gamma^{\prime})\right)}{C(\theta)^{2}}\mathrm{d}\Gamma^{\prime}$
@@
In this work, we use the following unbiased estimator for covariance. Given 2 random vectors $X,Y$, and samples of these vectors $\{(X^{(1)},Y^{(1)}),...(X^{(n)},Y^{(n)})\}$, our estimator is given by:
$\widehat{\operatorname{Cov}}(X,Y)=\frac{1}{N-1}\sum_{i=1}^{N}X^{(i)}{Y^{(i)}}^{\top}-\frac{1}{N(N-1)}\left(\sum_{j=1}^{N}X^{(j)}\right)\left(\sum_{k=1}^{N}{Y^{(k)}}^{\top}\right).$
@@
$=\frac{1}{N-1}\sum_{i=1}^{N}\mathbb{E}[X^{(i)}{Y^{(i)}}^{\top}]-\frac{1}{N(N-1)}\left(\sum_{i=1}^{N}\mathbb{E}[X^{(i)}{Y^{(i)}}^{\top}]+\sum_{j\neq k}\mathbb{E}[X^{(j)}{Y^{(k)}}^{\top}]\right)$
@@
Given the samples $\{\Gamma^{(1)},,...\Gamma^{(N)}\}$, we can then use the estimator above to get,
$\widehat{\operatorname{Cov}}\left(g(\Gamma),\nabla_{\theta}U_{\theta}(\Gamma)\right)$ $=\frac{1}{N-1}\sum_{i=1}^{N}g(\Gamma^{(i)})\nabla_{\theta}U_{\theta}(\Gamma^{(i)})^{\top}-\frac{1}{N(N-1)}\left(\sum_{j=1}^{N}g(\Gamma^{(j)})\right)\left(\sum_{k=1}^{N}\nabla_{\theta}U_{\theta}(\Gamma^{(k)})\right)^{\top}$

### v2

@@ 4.3 Ac-Ala3-NHMe Tetrapeptide
Figure 5: Ac-Ala3-NHMe tetrapeptide simulation with StABlE Training.(a) Fraction of unstable replicas as a function of simulation time for NequIP MLFFs. Applying StABlE Training yields a model which can simulate more MD replicas stably over time than a baseline trained only on energies and forces. (b) Distribution of interatomic distances ($h(r)$) from NequIP MLFF simulations. The StABlE-trained NequIP model much more closely recovers the true $h(r)$, while the $h(r)$ produced by the model trained only on QM reference data is noisy and inaccurate due to insufficient sampling time. (c) Ac-Ala3-NHMe structures sampled over epochs of a single learning phase of StABlE Training. There is a clear progression as very unstable configurations become stable.
Ac-Ala3-NHMe (chemical formula $C_{12}H_{22}N_{4}O_{4}$) is a tetrapeptide from the MD22 dataset Chmiela et al. (2022), a challenging benchmark consisting of reference energy and forces computed at the PBE+MBD Perdew et al. (1996); Tkatchenko et al. (2012) level of theory for considerably larger molecules than those in the MD17 dataset. Relative to MD17 molecules, Ac-Ala3-NHMe poses unique challenges for atomistic simulations due to its larger size and flexibility Kabylda et al. (2023). As a result, more expressive MLFFs—sometimes incorporating E(3) equivariance Geiger & Smidt (2022)—are required to accurately model the underlying potential energy surface. In our experiments, we use a NequIP Batzner et al. (2022) model due to its promising accuracy and data efficiency on challenging atomistic simulation tasks Kozinsky et al. (2023); Merchant et al. (2023). We pre-train a NequIP model on the QM energy and forces matching objective using 14,890 (or 25%) of available structures from the training dataset. After convergence, we begin StABlE Training with the global Boltzmann estimator, using 48 parallel replicas at a temperature of 500K. As with aspirin (Sections 4.1 - 4.2), we detect instability using the maximum bond length deviation criterion, and choose the distribution of interatomic distances, $h(r)$, as our training observable. The StABlE-trained NequIP model gives an improvement in stability relative to a baseline model trained only on QM reference data (Figure 5a). Additionally, the StABlE-trained model closely recovers the true $h(r)$, while the baseline MLFF produces a highly noisy $h(r)$. This is because of insufficient sampling, due to a lower stable simulation time (Figure 5b). The baseline MLFF also underestimates the latter portion of the $h(r)$ distribution, indicating poor modelling of long-range interactions. Figure 5c shows selected structures throughout the course of a single learning phase of StABlE Training, demonstrating a progression from very unstable to stable structures.
@@ 4.4 Water
Figure 6: All-Atom Water Simulation with StABlE Training. (a) Example unstable configuration arising in a GemNet-T simulation of water. Instabilities are highly localized and manifest as unphysical bond stretches or intermolecular collisions, motivating the use of our localized Boltzmann estimator (Section 3.1) during training (b) Fraction of unstable replicas as a function of simulation time for GemNet-T MLFFs. StABlE Training yields a significant stability improvement relative to training on only energy and forces data. (c) Convergence of the diffusivity coefficient over MLFF simulations. The StABlE-trained model gives a considerably more accurate final estimate of the diffusivity coefficient. (d - f) Element-conditioned (OH, OO, HH) radial distribution functions (RDF) produced by MLFF simulations. The StABlE-trained model more accurately captures long-range interactions.
As a final evaluation of StABlE Training, we consider liquid-phase water. Water has historically posed unique challenges for atomistic simulations due to the presence of long-range interactions, proton disorder, and nuclear quantum effects Cheng et al. (2019); Raza et al. (2011); Markland & Ceriotti (2018). Additionally, estimating transport properties, such as the diffusivity coefficient, requires long, continuous trajectories to minimize statistical error. We use the dataset produced in Fu et al. Fu et al. (2022), which contains reference all-atom simulations of 64 water molecules using the flexible Extended Simple Point Charge model Wu et al. (2006) at 300K and 1 atm. For this system, we consider a GemNet-T model Gasteiger et al. (2021). GemNet-T displays a failure mode in which unphysical configurations (e.g., bond stretching) first arise in highly localized "pockets" of 1-2 molecules (Figure 6a), which then gradually cascade to the rest of the simulation domain Fu et al. (2022). Such configurations do not noticeably affect global system observables such as the radial distribution function. Therefore, in this setting it is more appropriate to use our localized version of the Boltzmann gradient estimator (Equation 5) in order to guide the optimization process towards specific spatial domains of instability. After pre-training a GemNet-T model using a subset of 1,000 reference structures, we begin StABlE Training with the localized Boltzmann estimator, simulating 5 parallel replicas at a temperature of 300K. We use a minimum intermolecular distance criterion (Section A.5) to detect instabilities. To estimate gradients for optimization, from each global state we extract all possible local neighborhoods containing a single water molecule (Supplementary Section A.6). The most prevalent failure mode produced by GemNet-T in this system is unphysical bond stretching, which eventually leads to unphysical coordination structures. Motivated by this observation, we use the mean hydrogen-oxygen bond length as our training observable. We perform a single cycle of simulation and learning. We then evaluate the performance of the StABlE-Trained model by performing 5 parallel nanosecond (ns) MD simulations at 300K, starting from held-out initial conditions. We achieve significant improvements in stability and accuracy using our training approach. Our StABlE-trained model can simulate stably for 1 ns for all 5 initial conditions, while the baseline MLFF model trained only on QM reference data can only do so for 1 initial condition, and has a median stability of just 312 ps (Figure 6b). Compared to a reference value of $2.3\times 10^{-9}\frac{m^{2}}{s}$, the mean diffusivity produced by StABlE-trained model simulations is $2.4\times 10^{-9}\frac{m^{2}}{s}$, while the mean diffusivity produced by baseline model simulations is $1.3\times 10^{-9}\frac{m^{2}}{s}$ (Figure 6c). The long-range correlations in the element-conditioned RDFs (Figures 6d-f) are also captured more accurately by the StABlE-trained model. We reiterate that neither the element-conditioned RDFs nor the diffusivity coefficient were explicitly seen during StABlE Training. In Supplementary Section A.10, we demonstrate that StABlE Training leads to similar stability improvements when water simulations are conducted in the isothermal/isobaric (NPT) ensemble, instead of the canonical (NVT) ensemble.
5 Conclusion and Future Work
We have introduced StABlE Training, a strategy for training stable and accurate machine learning force fields. Our training procedure results in MLFFs which are significantly more stable in MD simulation, more accurately reproduce key simulation observables, (including those that were not explicitly trained on), exhibit better generalization to unseen temperatures, and have superior data efficiency relative to MLFFs trained only on QM data.
Key Takeaways.
StABlE Training can utilize both quantum-mechanical energies and forces and system observables to supervise MLFF training. We highlight that the reference observables need not be acquired from experimental measurements, and can be computed by averaging over existing QM datasets. This makes our approach applicable in realistic computational discovery scenarios in which experimental characterization is not always available for hypothetical systems. The stability, accuracy, and data efficiency gains brought by StABlE Training require no additional reference calculations or data. The procedure is thus self-contained and efficient, requiring minimal additional computational expense beyond a single iteration of conventional MLFF training (see Supplementary Section A.6 for details on training times). This suggests that StABlE Training may scale more gracefully to larger MLFFs and atomistic systems than existing active learning approaches, which require repeated QM reference calculations and retraining. StABlE Training can also be flexibly applied across atomistic systems due to its ability to handle a diverse set of failure modes arising in MD simulation, including both global and local instabilities due to our localized Boltzmann estimator. This flexibility extends to the choice of MLFF: the effectiveness of StABlE Training across the three diverse architectures considered in this work suggests that our approach will remain applicable as MLFF architectures continue to evolve.
@@ 5 Conclusion and Future Work
Unlike traditional QM learning, training MLFFs to match reference observables lacks convergence guarantees in the large data limit. Specifically, the observable-matching objective is under-constrained, as the mapping between a potential energy function and a sparse set of simulation observables is in general non-unique Noid (2013). Although we mitigate this problem by pre-training and regularizing StABlE Training with the conventional energy and forces loss function, we observe that stability improvements resulting from StABlE Training are accompanied by a minor increase in the energy and forces error on a held-out test set, indicating a misalignment between the observable-matching and QM objectives. Adjusting the strength of the QM regularization and the learning rate used for StABlE Training are two primary ways in which to navigate this tradeoff. We note, however, that our observed error increases are typically within the range of DFT error (Supplementary Section A.11). We also note that StABlE Training is currently incompatible with dynamical observables, due to the use of uncorrelated states to compute the Boltzmann Estimator. Overcoming this limitation is non-trivial, as it requires optimization over long paths, which is a significantly higher-dimensional problem with high-variance gradients Bolhuis et al. (2023).
@@
In this work, we took observables calculated from high-fidelity simulations as reference values. Many of the observables we considered, such as the radial distribution function, diffusivity coefficient, and equilibrium bond length, are also experimentally measurable. Future work could explore training with experimental observables from multiple thermodynamic conditions simultaneously to yield more generalizable and robust potentials. Future work could also explore incorporating additional observables into StABlE Training, particularly those which are dynamical. This could restrict the set of learnable functions and address the under-constrained nature of learning with observables. We also note that the reference datasets used to pre-train MLFFs in this work were sampled uniformly from high-fidelity simulations. More sophisticated sampling strategies Deringer et al. (2018); Karabin & Perez (2020); Yoo et al. (2021); Fonseca et al. (2021); Kulichenko et al. (2023); Qi et al. (2024) for diverse dataset generation, as well as during phase space exploration to find unstable regions, could be employed along with StABlE Training to achieve further stability and accuracy gains. Closely related techniques such as active learning are compatible with StABlE Training, and could be used in tandem to further improve MLFF stability and robustness. In this setting, system observables can serve as a cheap source of information with which to augment the more expensive supervision provided by ab-initio calculations.
Appendix A Appendix
@@ A.1 Derivation of Boltzmann Estimator
We provide a full derivation of our Boltzmann estimator, which we use to train our MLFF as part of StABlE Training. Consider a vector-valued observable $g(\Gamma)$ of a state $\Gamma$, and a reference value of the observable $g_{\text{ref}}$. Training a MLFF $U_{\theta}$ to match $g_{\text{ref}}$ requires minimizing the loss function,
@@
where $P_{\theta}(\Gamma)$ is the equilibrium distribution induced by the MLFF $U_{\theta}$. This requires computing the gradient, $\nabla_{\theta}\mathcal{L}_{\text{obs}}$, which can be decomposed via the chain rule as follows:
@@
We derive the $N$-sample estimator, presented in Equation 6 in the main text, of the Jacobian, ${\partial\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)}[g(\Gamma)]}/{\partial\theta}$. The estimator is repeated below for convenience.
$\mathcal{E}(\Gamma_{1},\ldots,\Gamma_{N})\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{N}{k_{B}T(N-1)}\left[\hat{\mathbb{E}}\left[g(\Gamma)\right]\hat{\mathbb{E}}\left[{\nabla_{\theta}U_{\theta}(\Gamma)}\right]^{\top}-\hat{\mathbb{E}}\left[{g(\Gamma)\nabla_{\theta}U_{\theta}(\Gamma)}^{\top}\right]\right],$
where $\hat{\mathbb{E}}$ denotes sample averages. This is an unbiased estimator, that is, $\mathbb{E}_{\Gamma_{1},\ldots,\Gamma_{N}\sim P_{\theta}(\Gamma)}[\mathcal{E}(\Gamma_{1},\ldots,\Gamma_{N})]=\frac{\partial\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)}[g(\Gamma)]}{\partial\theta}$.
@@
Given $P_{\theta}(\Gamma)\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{\exp\left(-\frac{1}{k_{B}T}\mathcal{H}_{\theta}(\Gamma)\right)}{C(\theta)}$, where $\mathcal{H}_{\theta}(\Gamma)=\sum_{i=1}^{N}\frac{p_{i}^{2}}{2m_{i}}+U_{\theta}(r)$, and $C(\theta)\stackrel{{\scriptstyle\text{def}}}{{=}}\int\exp(-\frac{1}{k_{B}T}\mathcal{H}_{\theta}(\Gamma^{\prime}))\mathop{}\!\mathrm{d}\Gamma^{\prime}$ is the partition function, we have,
@@
Expanding via the chain rule for Jacobians, and noting that $\nabla_{\theta}\mathcal{H}_{\theta}(\Gamma)=\nabla_{\theta}U_{\theta}(r)$ (we will write $U_{\theta}(\Gamma)$ for convenience) since the kinetic energy is independent of $\theta$, we get,
$\frac{\partial\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)}[g(\Gamma)]}{\partial\theta}$ $=\int g(\Gamma^{\prime})\frac{-\frac{1}{k_{B}T}\exp\left(-\frac{1}{k_{B}T}\mathcal{H}_{\theta}(\Gamma^{\prime})\right)\nabla_{\theta}U_{\theta}(\Gamma^{\prime})^{\top}C(\theta)-\nabla_{\theta}C(\theta)^{\top}\exp\left(-\frac{1}{k_{B}T}\mathcal{H}_{\theta}(\Gamma^{\prime})\right)}{C(\theta)^{2}}\mathrm{d}\Gamma^{\prime}$
@@
In this work, we use the following unbiased estimator for covariance. Given 2 random vectors $X,Y$, and samples of these vectors $\{(X_{1},Y_{1}),...(X_{n},Y_{n})\}$, our estimator is given by:
$\widehat{\operatorname{Cov}}(X,Y)=\frac{1}{N-1}\sum_{i=1}^{N}X_{i}{Y_{i}}^{\top}-\frac{1}{N(N-1)}\left(\sum_{j=1}^{N}X_{j}\right)\left(\sum_{k=1}^{N}{Y_{k}}^{\top}\right).$
@@
$\widehat{\operatorname{Cov}}(X,Y)$ $=\frac{1}{N-1}\sum_{i=1}^{N}\mathbb{E}[X_{i}{Y_{i}}^{\top}]-\frac{1}{N(N-1)}\left(\sum_{i=1}^{N}\mathbb{E}[X_{i}{Y_{i}}^{\top}]+\sum_{j\neq k}\mathbb{E}[X_{j}{Y_{k}}^{\top}]\right)$
@@
Given the samples $\{\Gamma_{1},,...\Gamma_{N}\}$, we can then use the estimator above to get,
$\widehat{\operatorname{Cov}}\left(g(\Gamma),\nabla_{\theta}U_{\theta}(\Gamma)\right)$ $=\frac{1}{N-1}\sum_{i=1}^{N}g(\Gamma_{i})\nabla_{\theta}U_{\theta}(\Gamma_{i})^{\top}-\frac{1}{N(N-1)}\left(\sum_{j=1}^{N}g(\Gamma_{j})\right)\left(\sum_{k=1}^{N}\nabla_{\theta}U_{\theta}(\Gamma_{k})\right)^{\top}$